OpenStudy (anonymous):

Who wants to help on parabolas? x^2 + 3y = 0 Graph how?

hero (hero):

Hint make an (x,y) table Start with (x,y) = (0,0)

hero (hero):

To get the next point, substitute x = 1 into the equation and solve for y: x^2 + 3y = 0 (1)^2 + 3y = 0 1 + 3y = 0 3y = -1 y = -1/3 So the next point is (1, -1/3)

hero (hero):

Keep going until you have enough points to plot an appropriate graph

hero (hero):

Solving for y will allow you to plot what the graph looks like on a calculator: x^2 + 3y = 0 3y = -x^2 y = -x^2/3

OpenStudy (anonymous):

If you don't have a graphing calculator, you can use geogebra, use Hero's last equation on the input line. http://www.geogebra.org/cms/

hero (hero):

I've attached what the graph looks like. You should be able to do the rest.

OpenStudy (anonymous):

THANK YOU ! !